# MoSA: Motion-Coherent Human Video Generation via Structure-Appearance Decoupling

### v2

@@ 1 Introduction
Existing video generation models (Kong et al., 2024; Yang et al., 2024b; Wan et al., 2025; Ma et al., 2025) often lack explicit guidance from human structural priors and are typically trained with noise reconstruction objectives in pixel space. Previous studies (Chefer et al., 2025; Jeong et al., 2024) have demonstrated that this paradigm leads to an overemphasis on appearance fidelity while neglecting human structural coherence, resulting in unrealistic human motion in the generated videos, as shown in Fig. 1(a). Since human appearance and motion convey different cues, they should adhere to different generation paradigms. This intuition leads to our MoSA, which generates human video through decoupling the structure and appearance generation. Specifically, as it is difficult to generate human videos with complex motions directly from the given text prompt, we first generate human motion structures conditioned on the text prompt. Given the sparse motion structures, MoSA subsequently synthesizes the visual appearance.
@@
Figure 1: Illustration of the motivation. (a) shows sampled frames from videos generated with the prompt “running”, where existing works (Genmo, 2024; Yang et al., 2024b) struggle to generate human videos with reasonable structures. (b) compares existing human video datasets (Wang et al., 2024d; Jafarian & Park, 2021) and our Movid, where existing datasets mostly focus on facial or upper-body regions, or consist of vertically oriented dance videos. More samples Movid are provided in Fig. 7 and supplementary materials.
@@
Besides the above methodology, this paper also contributes a novel human video dataset presenting complex human motions. Most existing human video datasets (Yu et al., 2023; Wang et al., 2024d; Li et al., 2024) primarily capture facial and upper-body movements with relatively simple movements, as illustrated in Fig. 1(b). Similarly, existing dance datasets (Jafarian & Park, 2021; Castro et al., 2018) exhibit restricted background diversity and motion complexity, which confines corresponding generation approaches (Hu, 2024; Hu et al., 2025; Zhu et al., 2024; Gan et al., 2025; Zhang et al., 2024b) to dance videos and often requires auxiliary pose inputs. The fact that most existing open-source human video datasets (Wang et al., 2024d; Li et al., 2024; Jafarian & Park, 2021; Yu et al., 2023) primarily focus on simple movements, makes models trained on such datasets struggle to generate realistic and physically plausible motions. We thus introduce MoVid, a novel dataset comprising 30K human motion videos exhibiting diverse action categories and complex motions.
@@ 2 Related Work
Human Motion Generation. Text-driven human motion generation aims to produce a sequence of human keypoints conditioned on a given text prompt, which can then be transformed into structural representations such as skeletons. Several existing approaches (Zhang et al., 2023; Chen et al., 2023; Tevet et al., 2022; Zhang et al., 2024a; Yuan et al., 2023; Meng et al., 2024b; Guo et al., 2024; Fan et al., 2025) leverage models trained on 3D motion-annotated datasets (Guo et al., 2022; Plappert et al., 2016; Lin et al., 2023; Zhang et al., 2025b) to generate 3D keypoint sequences. For example, MLD (Chen et al., 2023) extends latent diffusion models to support text-to-motion generation, while T2M-GPT (Zhang et al., 2023) employs a generative pre-trained transformer (Radford et al., 2018) as the backbone and utilizes a vector-quantized variational autoencoder (Van Den Oord et al., 2017) to encode and reconstruct keypoint features. In addition, several studies (Wang et al., 2025a; 2024c; Song et al., 2024) have explored directly generating 2D keypoints or skeleton sequences as representations of human motion.
@@ 3.1 Preliminary
Video Generation Model. Diffusion transformer (DiT) (Peebles & Xie, 2023) based generative models have attracted increasing attention due to their strong performance and scalability. Building on this (Yang et al., 2024b; Wan et al., 2025), we adopt DiT as the backbone and extend it to support motion-coherent human video generation.
@@ 3.3 Human-Aware Dynamic Control
Method FVD CLIPSIM
Subject
Consistency
Background
Consistency
Motion
Smoothness
Dynamic
Degree
Imaging
Quality
@@
Figure 3: Visual comparison with existing video generation models. For clarity, VideoCrafter2 (Chen et al., 2024) is denoted as VC2, and HunyuanVideo (Kong et al., 2024) is denoted as Hunyuan.
@@ 6 Conclusion
  • Details of the proposed 3D contact constraint in Sec. B.
@@
  • More comparative experiments in Sec. E, including comparisons with general video diffusion models, text-driven human video generation models, commercial video generation models, and human animation models.
  • More ablation studies in Sec. F, including details of the 2D structure generation model, effect of human-aware dynamic control modules, effect of dense tracking loss, effect of the contact constraint, necessity of the MoVid dataset, and effect of the selected camera poses.
@@
  • LLM usage declaration in Sec. I.
@@ Appendix B Details of the Proposed 3D Contact Constraint
To ensure physically plausible human–environment interactions, we introduce a 3D contact loss that penalizes unrealistic interpenetrations. The computation proceeds in four steps. First, each video frame is lifted into a 3D point cloud representation using a pretrained VGGT (Wang et al., 2025b) model. Formally, for the $f$-th frame, we obtain a point set $P_{f}={p_{i}\in\mathbb{R}^{3}}$. Second, human and scene points are separated based on 2D segmentation masks. Each 3D point $p_{i}$ is projected onto the image plane using the intrinsic matrix $K_{f}$ and extrinsic matrix $E_{f}$ predicted by the VGGT model. If the projected pixel $(u_{i},v_{i})$ lies within the segmented human region $M_{f}$, the point is assigned to the human set $H_{f}$; otherwise, it is assigned to the scene set $S_{f}$. Third, scene points from all frames are aggregated to reconstruct a mesh $M_{b}$ via convex hull estimation. This mesh is then converted into a signed distance function (SDF), denoted $SDF(\cdot)$, where negative values indicate locations inside the scene surface and positive values denote outside regions. Finally, the 3D contact loss is defined. If human points penetrate the scene ($SDF(h)<\tau$), the loss penalizes the penetration depth; otherwise, it encourages human points to remain close to the scene surface by minimizing their distance to the mesh. The resulting loss is formally expressed as:
@@ C.2 Dataset Statistics
Tab. 8 provides a detailed comparison with several representative real-world human video datasets, including CelebV-HQ (Zhu et al., 2022), CelebV-Text (Yu et al., 2023), TikTok (Jafarian & Park, 2021), UBC-Fashion (Zablotskaia et al., 2019), IDEA-400 (Lin et al., 2023) and HumanVid (Wang et al., 2024d). Most existing datasets, such as CelebV (Yu et al., 2023; Zhu et al., 2022), HumanVid (Wang et al., 2024d), and the recent OpenHumanVid (Li et al., 2024), predominantly focus on facial or upper-body motions. In addition, other datasets are often constrained by limited motion diversity or specific video formats. For example, TikTok (Jafarian & Park, 2021) and UBC-Fashion (Zablotskaia et al., 2019) primarily consist of vertically oriented videos with restricted motion ranges. Others, such as IDEA-400 (Lin et al., 2023), lack fine-grained and accurate textual annotations, which limits their applicability in text-conditioned generation tasks. To address these limitations, we introduce the MoVid, a high-quality whole-body human video dataset with millions of frames. Examples are provided in dataset_sample.zip of the provided supplementary materials, and are also shown in Fig. 7.
@@ Appendix D Implementation details of MoSA
We use CogVideoX-5B-T2V (Yang et al., 2024b) as the base model of the appearance generation branch. During training, we freeze the original weights and train only the structure generation blocks and HADC modules. We also present the results for the version using Wan 2.1 (Wan et al., 2025) as the base model in Tab. 2 and Fig. 4, following the same training settings as described above. For the proposed HADC modules, $\mathcal{P}^{k}$ comprises three linear layers interleaved with two activation functions (Hendrycks & Gimpel, 2016). $\mathcal{U}^{k}$ shares a similar architecture with $\mathcal{P}^{k}$, but includes an additional up-sampling layer followed by a 3D convolution layer (Tran et al., 2015) at the end. During training, the skeleton sequence is extracted from the input video $V$. We train and test at a resolution of 720×480, and train for 20,000 iterations on 4 NVIDIA A800 GPUs with a batch size of 16. The AdamW (Loshchilov & Hutter, 2017) optimizer is used with a learning rate of 1e-5. We set the loss weights $\lambda_{m}$, $\lambda_{track}$ and $\lambda_{cont}$ to 0.001, 0.01 and 10.0, respectively.
@@ E.4 Comparison with the commercial models of Kling and Seedance
Method FVD CLIPSIM
Subject
Consistency
Background
Consistency
Motion
Smoothness
Dynamic
Degree
Imaging
Quality
@@ Appendix H Image-to-Video Generation variant of MoSA
Appendix I LLM Usage Declaration

### v3

@@ 1 Introduction
Existing video generation models (Kong et al., 2024; Yang et al., 2024b; Wan et al., 2025) often lack explicit guidance from human structural priors and are typically trained with noise reconstruction objectives in pixel space. Previous studies (Chefer et al., 2025; Jeong et al., 2024) have demonstrated that this paradigm leads to an overemphasis on appearance fidelity while neglecting human structural coherence, resulting in unrealistic human motion in the generated videos, as shown in Fig. 1(a). Since human appearance and motion convey different cues, they should adhere to different generation paradigms. This intuition leads to our MoSA, which generates human video through decoupling the structure and appearance generation. Specifically, as it is difficult to generate human videos with complex motions directly from the given text prompt, we first generate human motion structures conditioned on the text prompt. Given the sparse motion structures, MoSA subsequently synthesizes the visual appearance.
@@
Figure 1: Illustration of the motivation. (a) shows sampled frames from videos generated with the prompt “running”, where existing works (Genmo, 2024; Yang et al., 2024b) struggle to generate human videos with reasonable structures. (b) compares existing human video datasets (Wang et al., 2024d; Jafarian and Park, 2021) and our Movid, where existing datasets mostly focus on facial or upper-body regions, or consist of vertically oriented dance videos. More samples Movid are provided in Fig. 7 and supplementary materials.
@@
Besides the above methodology, this paper also contributes a novel human video dataset presenting complex human motions. Most existing human video datasets (Yu et al., 2023; Wang et al., 2024d; Li et al., 2024) primarily capture facial and upper-body movements with relatively simple movements, as illustrated in Fig. 1(b). Similarly, existing dance datasets (Jafarian and Park, 2021) exhibit restricted background diversity and motion complexity, which confines corresponding generation approaches (Hu, 2024; Hu et al., 2025; Zhu et al., 2024; Gan et al., 2025; Zhang et al., 2024b) to dance videos and often requires auxiliary pose inputs. The fact that most existing open-source human video datasets (Wang et al., 2024d; Li et al., 2024; Jafarian and Park, 2021; Yu et al., 2023) primarily focus on simple movements, makes models trained on such datasets struggle to generate realistic and physically plausible motions. We thus introduce MoVid, a novel dataset comprising 30K human motion videos exhibiting diverse action categories and complex motions.
@@ 2 Related Work
Human Motion Generation. Text-driven human motion generation aims to produce a sequence of human keypoints conditioned on a given text prompt, which can then be transformed into structural representations such as skeletons. Several existing approaches (Zhang et al., 2023a; Chen et al., 2023; Tevet et al., 2022; Zhang et al., 2024a; Yuan et al., 2023; Meng et al., 2024b; Guo et al., 2024; Fan et al., 2025) leverage models trained on 3D motion-annotated datasets (Guo et al., 2022; Plappert et al., 2016; Lin et al., 2023; Zhang et al., 2025b) to generate 3D keypoint sequences. For example, MLD (Chen et al., 2023) extends latent diffusion models to support text-to-motion generation, while T2M-GPT (Zhang et al., 2023a) employs a generative pre-trained transformer (Radford et al., 2018) as the backbone and utilizes a vector-quantized variational autoencoder (Van Den Oord et al., 2017) to encode and reconstruct keypoint features. In addition, several studies (Wang et al., 2025a; 2024c; Song et al., 2024) have explored directly generating 2D keypoints or skeleton sequences as representations of human motion.
@@ 3.1 Preliminary
Video Generation Model. Diffusion transformer (DiT) (Peebles and Xie, 2023) based generative models have attracted increasing attention due to their strong performance and scalability. Building on this (Yang et al., 2024b; Wan et al., 2025), we adopt DiT as the backbone and extend it to support motion-coherent human video generation.
@@ 3.3 Human-Aware Dynamic Control
Method FVD CLIPSIM Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Imaging Quality
@@
Figure 3: Visual comparison with existing video generation models. For clarity, VideoCrafter2 (Chen et al., 2024) is denoted as VC2, and HunyuanVideo (Kong et al., 2024) is denoted as Hunyuan. It demonstrates that our method is capable of generating realistic human motion with with reasonable structures, while simultaneously preserving high fidelity in video appearance.
@@ 6 Conclusion
7 discussions and future work
Generating complex inter-person contact and fine-grained hand interactions has long been a challenging problem for video generation models. Built upon a structure-appearance disentanglement paradigm, our method substantially improves the quality of multi-person interactions and fine-grained motion synthesis. As shown in Fig. 17, the girl’s bowstring-pulling action in the second row and the two-person jumping and ball-contesting scene in the third row demonstrate clear enhancements.
Nevertheless, we found that generating highly intricate hand motions remains a challenge, where current methods may still produce artifacts such as distorted or blurred finger structures. However, our structure-appearance disentanglement paradigm is naturally compatible with incorporating denser structural cues, such as hand keypoints. This inherent compatibility offers a promising pathway to enhance the realism of finger-level actions by integrating such detailed guidance.
Our further investigation reveals that the primary obstacle lies in the training data. Existing motion datasets used for training 3D structure transformers typically include only SMPL body joints. Consequently, incorporating hand keypoints would necessitate augmenting these datasets with additional 3D annotations for hand joints to enable effective model training. We believe this is a very worthwhile direction to explore and will pursue it in our future work, with extending the proposed MoVid dataset.
@@
Acknowledgments
This work was supported in part by under Grant 2023-JCJQ-LA-001-088, in part by the Natural Science Foundation of China under Grant U20B2052, Grant 61936011, in part by under 2025ZD1601300, in part by the Okawa Foundation Research Award, in part by the Ant Group Research Fund, and in part by the Kunpeng&Ascend Center of Excellence, Peking University.
@@ Acknowledgments
  • Details of the proposed 3D contact constraint in Sec. B, including its differential process.
@@
  • More comparative experiments in Sec. E, including more comparisons with general video diffusion models, text-driven human video generation models, commercial video generation models, human animation models, and generation results with occlusion situations.
  • More ablation studies in Sec. F, including details of the 2D structure generation model, final visualization of the HADC module outputs, effect of human-aware dynamic control modules, effect of dense tracking loss, effect of the contact constraint, necessity of the MoVid dataset, effect of the selected camera poses, and the effect of the data from different views in our dataset..
@@
  • Evaluation on more motion-centric metrics in Sec. I.
  • LLM usage declaration in Sec. J.
@@ Appendix B Details of the Proposed 3D Contact Constraint
To ensure physically plausible human–environment interactions, we introduce a 3D contact loss that penalizes unrealistic interpenetrations. The computation proceeds in four steps. First, each video frame is lifted into a 3D point cloud representation using a pretrained VGGT (Wang et al., 2025c) model. Formally, for the $f$-th frame, we obtain a point set $P_{f}={p_{i}\in\mathbb{R}^{3}}$. Second, human and scene points are separated based on 2D segmentation masks. Each 3D point $p_{i}$ is projected onto the image plane using the intrinsic matrix $K_{f}$ and extrinsic matrix $E_{f}$ predicted by the VGGT model. If the projected pixel $(u_{i},v_{i})$ lies within the segmented human region $M_{f}$, the point is assigned to the human set $H_{f}$; otherwise, it is assigned to the scene set $S_{f}$. Third, scene points from all frames are aggregated to reconstruct a mesh $M_{b}$ via convex hull estimation. This mesh is then converted into a signed distance function (SDF), denoted $SDF(\cdot)$, where negative values indicate locations inside the scene surface and positive values denote outside regions. Finally, the 3D contact loss is defined. If human points penetrate the scene ($SDF(h)<\tau$), the loss penalizes the penetration depth; otherwise, it encourages human points to remain close to the scene surface by minimizing their distance to the mesh. The resulting loss is formally expressed as:
@@
Differential Process. For the f-th frame, the 3D contact loss is defined as Eq. 10.
The propagation path is as follows, where $h$ denotes a human point, $V^{\prime}$ denotes the generated video, $\theta$ denotes the model parameters.
$\mathcal{L}_{cont}^{f}\rightarrow\text{SDF}(h)\rightarrow h\rightarrow V^{\prime}\rightarrow\theta$ (11)
(1) The process begins with the loss function $\mathcal{L}_{cont}^{f}$ itself. Since $\tau$ is set to 0, for a human point $h$, the loss form can be simplified to:
$\mathcal{L}_{cont}^{f,h}=\max(0,-\text{SDF}(h))$ (12)
(2) The partial derivative of this loss with respect to the SDF value is non-zero only in the case of interpenetration:
$\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial\text{SDF}(h)}=\begin{cases}-1,&\text{if }\text{SDF}(h)<0,\\ 0,&\text{otherwise}.\end{cases}$ (13)
(3) The next step is to propagate the gradient from the SDF value to the 3D coordinates $(x,y,z)$ of the point $h$. The gradient of a scalar field like the SDF with respect to a point’s coordinates is a vector:
$\nabla_{h}\text{SDF}=\frac{\partial\text{SDF}(h)}{\partial h}$ (14)
This gradient vector provides the most efficient direction to displace the point $h$ in order to move it towards the exterior of the scene mesh, thereby resolving the penetration.
Using the chain rule, we combine the gradients from the previous two steps to compute the gradient of the loss with respect to the 3D coordinates of the human point $h$:
$\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial h}=\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial\text{SDF}(h)}\cdot\frac{\partial\text{SDF}(h)}{\partial h}$ (15)
(4) Then, the gradient $\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial h}$ can be backpropagated through the layers of VGGT [2] to obtain the gradient with respect to its input, i.e. the generated video frame $V^{\prime}$:
$\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial V^{\prime}}=\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial h}\cdot\frac{\partial h}{\partial V^{\prime}}$ (16)
(5) Finally, the generated video frame $V^{\prime}$ is the output of our model, $V^{\prime}=\text{MoSA}(\text{inputs};\theta)$. With the gradient $\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial V^{\prime}}$ computed, we can continue the backpropagation through the MoSA network to obtain the gradients with respect to all of its learnable parameters $\theta$:
$\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial\theta}=\frac{\partial\mathcal{L}_{cont}^{f,h}}{\partial V^{\prime}}\cdot\frac{\partial V^{\prime}}{\partial\theta}$ (17)
These final gradients are then used by the optimizer to update the model’s parameters.
@@ C.2 Dataset Statistics
Tab. 8 provides a detailed comparison with several representative real-world human video datasets, including CelebV-HQ (Zhu et al., 2022), CelebV-Text (Yu et al., 2023), TikTok (Jafarian and Park, 2021), UBC-Fashion (Zablotskaia et al., 2019), IDEA-400 (Lin et al., 2023) and HumanVid (Wang et al., 2024d). Most existing datasets, such as CelebV (Yu et al., 2023; Zhu et al., 2022), HumanVid (Wang et al., 2024d), and the recent OpenHumanVid (Li et al., 2024), predominantly focus on facial or upper-body motions. In addition, other datasets are often constrained by limited motion diversity or specific video formats. For example, TikTok (Jafarian and Park, 2021) and UBC-Fashion (Zablotskaia et al., 2019) primarily consist of vertically oriented videos with restricted motion ranges. Others, such as IDEA-400 (Lin et al., 2023), lack fine-grained and accurate textual annotations, which limits their applicability in text-conditioned generation tasks. To address these limitations, we introduce the MoVid, a high-quality whole-body human video dataset with millions of frames. Examples are provided in dataset_sample.zip of the provided supplementary materials, and are also shown in Fig. 7.
@@ Appendix D Implementation details of MoSA
We use CogVideoX-5B-T2V (Yang et al., 2024b) as the base model of the appearance generation branch. During training, we freeze the original weights and train only the structure generation blocks and HADC modules. We also present the results for the version using Wan 2.1 (Wan et al., 2025) as the base model in Tab. 2 and Fig. 4, following the same training settings as described above. For the proposed HADC modules, $\mathcal{P}^{k}$ comprises three linear layers interleaved with two activation functions (Hendrycks and Gimpel, 2016; Wang et al., 2025b; Zhang et al., 2023b). $\mathcal{U}^{k}$ shares a similar architecture with $\mathcal{P}^{k}$, but includes an additional up-sampling layer followed by a 3D convolution layer (Tran et al., 2015) at the end. During training, the skeleton sequence is extracted from the input video $V$. We train and test at a resolution of 720×480, and train for 20,000 iterations on 4 NVIDIA A800 GPUs with a batch size of 16. The AdamW (Loshchilov and Hutter, 2017) optimizer is used with a learning rate of 1e-5. We set the loss weights $\lambda_{m}$, $\lambda_{track}$ and $\lambda_{cont}$ to 0.001, 0.01 and 10.0, respectively.
@@ E.4 Comparison with the commercial models of Kling and Seedance
Method FVD CLIPSIM Subject Consistency Background Consistency Motion Smoothness Dynamic Degree Imaging Quality
@@ E.5 Comparison with Human Animation Methods
E.6 generation results with occlusion situations
The core role of the HADC modules is to enhance the model’s fine-grained control over human motion. During training, we introduce a mask loss $L_{m}$ calculated from the ground-truth human masks. This process guides the model to learn how to effectively propagate the sparse structural guidance $s$ to the corresponding **visible regions** of the human body in the image. Through this supervision, the model learns the association between the skeletal structure and the actual visible body parts.
Consequently, this mechanism naturally handles scenarios with partial occlusions. During training, if a part of the human body is occluded by an object (e.g., a leg hidden behind a table), the corresponding ground-truth human mask for that area will be absent. The model learns that even when the structural guidance $s$ indicates that a leg should be present, it should avoid generating the leg’s appearance in regions where it is actually occluded. In other words, the HADC module learns to render the visual appearance only in the **non-occluded areas** corresponding to the structural guidance. This capability is demonstrated on the left side of Fig. 5 in our paper, where our model properly handles a person being occluded by objects.
To further substantiate the effectiveness of our approach, we add additional visual results in Fig. 18 in the appendix. These results showcase more diverse occlusion scenarios and clearly demonstrate that MoSA can generate high-quality and physically plausible videos even when partial occlusions are present.
@@ Appendix H Image-to-Video Generation variant of MoSA
Appendix I evaluation on more motion-centric metrics
We design more motion-centric metrics and add these evaluation metrics to the ablation study of the structure-appearance decoupling paradigm, as shown in Table 14. Furthermore, we use these metrics to compare existing T2V models with our approach, and the corresponding results are presented in Table 15. The Pose Confidence metric is computed by applying the DWPose [1] model to extract human keypoints from the generated videos and averaging their confidence scores. Higher confidence indicates more accurate pose estimation, which further reflects the plausibility of the generated human poses, avoiding structural failures such as missing limbs. Temporal Smoothness is derived by calculating the mean jerk of tracking points within human regions, where lower values correspond to smoother motion. The Contact-Violation metric is evaluated using Qwen3-VL, a powerful multimodal model equipped with extensive world knowledge. It identifies physically implausible artifacts such as interpenetrations in the generated videos and computes the corresponding violation rate. Action Recognizability is also assessed using Qwen3-VL, which determines whether the generated actions faithfully match those specified in the text prompt, such as the text prompt “walks to the blue trampoline and jumps” in Fig. 3. Motion Plausibility and Action Correctness are the evaluation criteria used in the user study.
Moreover, we have conducted a user study, as reported in Table 9 of the paper, to assess the motion quality of the generated videos. To more comprehensively demonstrate the advantages of our approach, we further evaluate the motion plausibility and action correctness of the generated videos, with the corresponding results presented in Table 16. These motion-centric metrics provide additional evidence that our method exhibits clear superiority in generating high-quality human motion.
Table 14: Effect of the decoupling paradigm, evaluating on more motion-centric metrics.
Structure Branch Pose Confidence Temporal Smoothness Contact- Violation Action Recognizability Motion Plausibility Action Correctness
0.885 0.122 2.04% 96.85% 5.09% 8.24%
2D Structure Generation 0.896 0.114 1.60% 98.42% 15.13% 15.60%
Ours 0.911 0.103 1.08% 99.45% 79.78% 76.16%
Table 15: Quantitative comparison with more motion-centric metrics.
Method Pose Confidence $\uparrow$ Temporal Smoothness $\downarrow$ Contact-Violation $\downarrow$ Action Recognizability $\uparrow$
VideoCrafter2 0.831 0.155 4.06% 89.70%
LaVie 0.865 0.138 3.25% 92.14%
Mochi 1 0.886 0.121 2.04% 96.75%
CogVideoX 0.882 0.124 2.17% 96.47%
HunyuanVideo 0.894 0.119 1.89% 97.28%
Wan2.1 0.893 0.117 1.62% 97.56%
Ours 0.911 0.103 1.08% 99.45%
Table 16: User study with more motion-aware metrics.
Method Motion Plausibility $\uparrow$ Action Correctness $\uparrow$
VideoCrafter2 0.85% 2.04%
LaVie 3.37% 2.14%
Mochi 1 12.18% 11.59%
CogVideoX 12.36% 14.82%
HunyuanVideo 16.45% 16.99%
Wan2.1 17.47% 17.14%
Ours 37.32% 35.28%
Table 17: Effect of the data from different views in our dataset.
Model FVD $\downarrow$ CLIPSIM $\uparrow$
w/o side-view data 1172 0.2959
w/o back-view data 1165 0.2963
Ours 1096 0.3031
Table 18: Ablation study on the human animation task.
Model SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$ FID $\downarrow$
w/o Struc. branch 0.677 20.603 0.266 682.1 42.74
w/o HADC 0.682 20.716 0.247 627.9 39.55
w/o Tracking loss 0.688 20.785 0.218 594.1 36.64
Ours 0.691 20.802 0.209 568.3 34.76
Appendix J LLM Usage Declaration
@@
Figure 18: Visual results in which the human body is partially occluded.